A couple's last child recently started high school. The couple would like to travel around the world once the child graduates from high school in 3.00 years. The couple currently has $38,698.00 in a money market fund that earns 3.12% APR with monthly compounding. How much will this account be worth when they are ready to travel the world?
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=38,698*(1+0.0312/12)^(12*3)

=38,698*1.09798703

=$42489.90(Approx).